OpenStudy (anonymous):

A 0.600 m long pendulum is used to determine the acceleration due to gravity on a distant planet. If 20 complete oscillations are completed in 35.5 s, the acceleration would be

OpenStudy (anonymous):

The period of a pendulum is \[T = 2\pi \sqrt {\frac {L}{g}}\]

OpenStudy (kropot72):

Note that the equation above is for a simple idealized body comprising a point mass suspended by a light inelastic cord. Under these conditions the equation applies only for small displacements.